Good morning!
Please help me with this. Solve the equation and then your the answer.
2(y+8)=3(y-6)

Answer by ankor@dixie-net.com(22740)

2(y+8) = 3(y-6)
:
2y + 16 = 3y - 18
:
16 + 18 = 3y - 2y
:
34 = y